The D121 and Z- value for a thermophilic spore in whole milk were experimentally determined as 30 s and 10.5oC respectively. Calculate D150 (min) for the spore.
1
Expert's answer
2020-07-22T06:03:30-0400

Here given decimal reduction time D at 121 oC(D121) and the value z for a thermophilic spore =30 s and 10.5oC.

we know that

"log(\frac{D_1}{D_2})=\frac{T_1 -T_2}{z}"

"log(\frac{30}{D})=\frac{121 -150}{10.5}"


"\frac{ 30}{D}=" antilog(-2.7619)


D=0.00087 min